OpenStudy (freckles):

I have a really dumb question about direction fields. I actually just started to try to learn about them.

OpenStudy (freckles):

\[\frac{dy}{dx}=x+y\] pretend we have this

OpenStudy (chosenmatt):

okay

OpenStudy (chosenmatt):

um....

OpenStudy (freckles):

\[0=x+y \text{ means } y=-x \] so for that one little part we have  that is suppose to be some horizontal lines there

OpenStudy (chosenmatt):

*reporting*

OpenStudy (chosenmatt):

oops wrong tag

OpenStudy (freckles):

i lost my train of thought

OpenStudy (freckles):

it had something to do at (0,0)

OpenStudy (freckles):

ok well whatever...  like when we have slope C is when y=-x+C